# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

### v1

@@ Abstract
Path-wise observables—functionals of stochastic trajectories—are at the heart of time-average statistical mechanics and are central to thermodynamic inequalities such as uncertainty relations, speed limits, and correlation-bounds. They provide a means of thermodynamic inference in the typical situation, when not all dissipative degrees of freedom in a system are experimentally accessible. So far, theories focusing on path-wise observables have been developing in two major directions, diffusion processes and Markov-jump dynamics, in a virtually disjoint manner. Moreover, even the respective results for diffusion and jump dynamics were derived with a patchwork of different approaches that are predominantly indirect. Stochastic calculus was recently shown to provide a direct approach to path-wise observables of diffusion processes, while a corresponding framework for jump dynamics remained elusive. In our work we develop, in an exact parallelism with continuous-space diffusion, a complete stochastic calculus for path-wise observables of Markov-jump processes. We formulate a “Langevin equation” for jump processes, define general path-wise observables, and establish their covariation structure, whereby we fully account for transients and time-inhomogeneous dynamics. We prove the known kinds of thermodynamic inequalities in their most general form and discus saturation conditions. We determine the response of path-wise observables to general (incl. thermal) perturbations and carry out the continuum limit to achieve the complete unification of diffusion and jump dynamics. Our results open new avenues in the direction of discrete-state analogs of generative diffusion models and the learning of stochastic thermodynamics from fluctuating trajectories.
@@ I Introduction
The manuscript is structured as follows. In Sec. II we first provide a summary of our main results. Next, in Sec. III we develop the stochastic calculus for jump processes, formalize general pathwise observables and establish their complete covariation structure. In Sec. IV we use the stochastic calculus to prove the (quite complete) list of thermodynamic inequalities in their most general form and apply them to thermodynamic inference on a selection of biophysical model systems. Following this, we show in Sec. V how the stochastic calculus can be applied to derive results for the response to general perturbations for discrete-state systems and how this relates to the fluctuation-dissipation theorem in equilibrium. In Sec. VI we take the continuum limit to show the equivalence of stochastic calculi for continuous- and discrete-state systems and in Sec. VII we extend the results to time-inhomogeneous dynamics. We conclude in Sec. VIII with an outlook. The mathematical proofs of the noise-time correlation Lemma and increment correlations, and other technical details are given in the Appendix A.
@@ II Summary of Main Results
for the $d$-dimensional position $\bm{x}_{\tau}$ in a drift field $\bm{F}(\bm{x}_{\tau})$ subject to thermal noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$, where $\mathrm{d}\bm{W}_{\tau}$ are increments of the $d$-dimensional Wiener process (i.e., $d$-dimensional “white noise”). In Sec. VII we will also allow for time-inhomogeneous (temporally driven) dynamics, while for clarity we will develop the results within the time-homogeneous setting.
@@
where the displacements have elements $(\mathrm{d}\bm{n}(\tau))_{xy}=\mathrm{d}n_{xy}(\tau)$ describing jumps from state $x$ to $y$ in $\tau,\tau+\mathrm{d}\tau]$. The matrix $(\bm{R}(x_{\tau},\tau))_{xy}\mathrm{d}\tau=\mathds{1}_{x_{\tau}=x}r_{xy}\mathrm{d}\tau$ encodes the expected (average) drift in $\tau,\tau+\mathrm{d}\tau]$ with time-independent transition rate $r_{xy}$ (following from the corresponding master equation with generator $(\bm{L})_{xy}=-\delta_{xy}\sum_{y\neq x}r_{xy}+(1-\delta_{xy})r_{yx}$), and the noise is encoded in the time-inhomogeneous shifted Poisson process $(\mathrm{d}\bm{\varepsilon}(\tau))_{xy}=\mathrm{d}\varepsilon_{xy}(\tau)$ [10, 134, 61, 60]. This noise has zero mean $\langle\mathrm{d}\varepsilon_{xy}(\tau)\rangle=0$ and the “white noise” property $\langle\mathrm{d}\varepsilon_{xy}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle_{x_{\tau}=k}=\delta_{ix}\delta_{jy}\delta_{ik}\delta(\tau-\tau^{\prime})r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$, where the expectation is conditioned on $x_{\tau}=k$.
We prove the central time-noise correlation lemma with $\mathrm{d}\tau_{i}(\tau)=\mathds{1}_{x_{\tau}=i}\mathrm{d}\tau$,
@@
$=\delta_{xk}\mathds{1}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau),$
@@
where $P(x,\tau^{\prime}|i,\tau)$ is the transition probability (i.e., the “propagator”) from state $x$ at $\tau$ to state $i$ at $\tau\leq\tau^{\prime}$ and $p_{i}(\tau)$ is the probability to find the system in state $i$ at time $\tau$. Note that the averages are “pinned”, e.g., $\langle\cdot\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}$ is the average conditioned on $x_{\tau}=i$} and $x_{\tau^{\prime}}=x$} at times $\tau\leq\tau^{\prime}$. The above correlations were shown in [60] using a different approach.
@@
$J_{t}^{\mathrm{c}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\circ\mathrm{d}\bm{x}_{\tau},$
@@
for continuous-space and MJP, respectively, where the former is a Stratonovich integral over some vector-valued $\bm{U}(\bm{x}_{\tau},\tau)$. For MJP the transitions are weighted by a time-dependent antisymmetric transition weight $\bm{\kappa}(\tau)=-\bm{\kappa}(\tau)$. Hence, we identify Stratonovich integrals for MJP as integrals over traces of a two-point function and the jump increments. The above currents naturally decompose into a stochastic and “usual” integral [61], $J^{(\mathrm{c})}_{t}=J_{t}^{\mathrm{(c),I}}+J_{t}^{\mathrm{(c),II}}$, where the stochastic part is
$J_{t}^{\mathrm{c,I}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$
@@
respectively, while the ”usual” part reads $J_{t}^{\mathrm{(c),II}}=\int_{0}^{t}\mathcal{U}_{\tau}^{(c)}\mathrm{d}\tau$ where $\mathcal{U}_{\tau}^{c}\equiv\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{F}(\bm{x}_{\tau})+\nabla\cdot\left[\bm{D}\bm{U}(\bm{x}_{\tau},\tau)\right]$ and $\mathcal{U}_{\tau}\equiv\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\bm{R}(x_{\tau})]$. Thus, we identify in the second line of Eq. (6) the Ito-type integral for observables of MJP by analogy to diffusions.
@@
$\rho_{t}$ $=\int_{\tau=0}^{\tau=t}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau),$ (7)
for continuous and discrete dynamics, respectively, where $V(\bm{x}_{\tau})$ and $\bm{V}(\tau)=(V_{1}(\tau),V_{2}(\tau),\dots)^{T}$ are scalar- and vector-valued functions, respectively.
@@
$\hat{{I}}^{t}_{U^{\alpha},U^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}U_{xy}^{\alpha}\sum_{i,j}U^{\beta}_{ij}[\cdot],$ (8)
for continuous and discrete dynamics respectively. The operator in the first line of Eq (8) is already known [96, 125]. For $X_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c)},k},\rho_{t}^{\mathrm{(c)},k}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t}^{\mathrm{(c)},l},\rho_{t}^{\mathrm{(c)},l}\}$, $k,l\in\{1,2\}$, we find
@@
where $X_{t}^{\mathrm{(c)}},Y^{\mathrm{(c)}}_{t}$ are the appropriate pathwise observables and $U^{X},U^{Y}$ “window functions” that define them ($\bm{\kappa}$ in the case of MJP). Moreover, $m\in\{1,2,3\}$ indicates density-density ($m=1$), current-density ($m=2$), and current-current ($m=3$) covariances. For the density covariance ($m=1$), $X_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},k}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t}^{\mathrm{(c)},l}$, we get
${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle\mathds{1}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (10)
@@
$\mathrm{cov}({\rho}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right]$
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right],$
$\mathrm{cov}(J_{t}^{k},{J}_{t}^{l})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (13)
where $V^{\alpha/\beta}\delta$ is a shorthand notation for $V^{\alpha/\beta}_{i}\delta_{ij}$, $J^{\mathrm{s}}_{xy}$ is the local steady-state current between $x$ and $y$, and $\hat{j}_{xy}^{\ddagger}$ is the so-called dual-reverse current operator (reversed steady-state current) [133, 96, 125] which for MJP reads $\hat{j}_{xy}=r_{xy}\to\hat{j}_{xy}^{\ddagger}=r_{yx}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$ [6]. The results for continuous-space diffusion are given in Refs. [96, 125]. For stationary initial conditions and with time-independent state function and transition weights, Eqs. (13) have the form of generalized Green-Kubo-like relations, in the sense that they relate (co)variances of observables to time integrals of (generalized) correlation functions of density- and current-observables (see Refs. [133, 96] for a discussion of diffusion processes). Notably, Eqs. (13) have a striking similarity to the continuous space expressions [96, 133]. Later we show how to to generalize Eqs. (13) to include time-dependent transition weights and state functions, as well as transient dynamics.
@@
where $c(t):\mathds{R}\to\mathds{R}$ is a deterministic weight function for the density and $\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}$ is the discrete (continuous) space entropy production. In analogy with [61] we introduced the discrete space modified time-integrated density and current $\tilde{\rho}_{t}=\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau)$ and $\tilde{J}_{t}=\int_{\tau=0}^{\tau=t}\tau\mathrm{Tr}\left[(\partial_{\tau}\bm{\kappa}(\tau))^{T}\mathrm{d}\bm{n}(\tau)\right]$ that account for the dependence of $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$ on $\tau$.
Setting $c(t)=0$ we recover the transient TUR. The CTUR (in contrast to the TUR) allows to turn the Cauchy-Schwarz inequality into an equality for all times given a suitable choice of $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$. This choice, however, requires explicit and detailed information about the system and is therefore not applicable in most realistic applications. We can nevertheless optimize Eq. (14) with respect to $c(t)$ for given $\bm{V}(\tau)$ and $\bm{\kappa}(\tau)$ to find

### v2

@@ Abstract
Path-wise observables—functionals of stochastic trajectories—are at the heart of time-average statistical mechanics and are central to thermodynamic inequalities such as uncertainty relations, speed limits, and correlation-bounds. They provide a means of thermodynamic inference in the typical situation, when not all dissipative degrees of freedom in a system are experimentally accessible. So far, theories focusing on path-wise observables have been developing in two major directions, diffusion processes and Markov-jump dynamics, in a virtually disjoint manner. Moreover, even the respective results for diffusion and jump dynamics were derived with a patchwork of different approaches that are predominantly indirect. Stochastic calculus was recently shown to provide a direct approach to path-wise observables of diffusion processes, while a corresponding framework for jump dynamics remained elusive. In our work we develop, in an exact parallelism with continuous-space diffusion, a complete stochastic calculus for path-wise observables of Markov-jump processes. We formulate a “Langevin equation” for jump processes, define general path-wise observables, and establish their covariation structure, whereby we fully account for transients and time-inhomogeneous dynamics. We prove the known kinds of thermodynamic inequalities in their most general form and discus saturation conditions. We determine the response of path-wise observables to general (incl. thermal) perturbations and introduce a corresponding response-function formalism. We carry out the continuum limit to achieve the complete unification of diffusion and jump dynamics. In addition, we connect the framework to quantum unraveling and the Belavkin equation for open quantum systems, associating quantum and classical descriptions of thermal systems. Our results open new avenues in the direction of discrete-state analogs of generative diffusion models and the learning of stochastic thermodynamics from fluctuating trajectories.
@@ I Introduction
The manuscript is structured as follows. In Sec. II we first provide a summary of our main results. Next, in Sec. III we develop the stochastic calculus for jump processes, formalize general pathwise observables and establish their complete covariation structure. In Sec. IV we use the stochastic calculus to prove the (quite complete) list of thermodynamic inequalities in their most general form and apply them to thermodynamic inference on a selection of biophysical model systems. Following this, we show in Sec. V how the stochastic calculus can be applied to derive results for the response to general perturbations for discrete-state systems, establish the corresponding response-function formalism, and show how this relates to the fluctuation-dissipation theorem in equilibrium. In Sec. VI we take the continuum limit to show the equivalence of stochastic calculi for continuous- and discrete-state systems and in Sec. VII we extend the results to time-inhomogeneous dynamics. Subsequently, in Sec. VIII, we discuss how the classical stochastic-calculus approach is connected to quantum unraveling and the (discrete) Belavkin equation and state its classical analogue. We conclude in Sec. IX with an outlook. The mathematical proofs of the noise-time correlation Lemma and increment correlations, and other technical details are given in the Appendix A.
@@ II Summary of Main Results
for the $d$-dimensional position $\bm{x}_{\tau}$ in a drift field $\bm{F}(\bm{x}_{\tau})$ subject to thermal noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$, where $\mathrm{d}\bm{W}_{\tau}$ are increments of the $d$-dimensional Wiener process (i.e., $d$-dimensional “white noise”), with $\langle\mathrm{d}\bm{W}_{\tau}\rangle=0$ and $\langle(\mathrm{d}\bm{W}_{\tau})_{i}(\mathrm{d}\bm{W}_{\tau^{\prime}})_{j}\rangle=\delta_{ij}\delta(\tau-\tau^{\prime})\mathrm{d}\tau\mathrm{d}\tau^{\prime}$, and $\bm{\sigma}$ is the noise amplitude assumed to have full rank (nonzero determinant). In Sec. VII we will also allow for time-inhomogeneous (temporally driven) dynamics, while for clarity we will develop the results within the time-homogeneous setting.
@@
where the displacements have elements $(\mathrm{d}\bm{n}(\tau))_{xy}=\mathrm{d}n_{xy}(\tau)$ describing jumps from state $x$ to $y$ in $\tau,\tau+\mathrm{d}\tau]$. The matrix $(\bm{R}(x_{\tau},\tau))_{xy}\mathrm{d}\tau={\color[rgb]{0,0,0}\delta_{x_{\tau}x}}r_{xy}\mathrm{d}\tau$ encodes the expected (average) drift in $\tau,\tau+\mathrm{d}\tau]$ with time-independent transition rate $r_{xy}$ (following from the corresponding master equation with generator $(\bm{L})_{xy}=-\delta_{xy}\sum_{y\neq x}r_{xy}+(1-\delta_{xy})r_{yx}$), and the noise is encoded in the time-inhomogeneous shifted Poisson process $(\mathrm{d}\bm{\varepsilon}(\tau))_{xy}=\mathrm{d}\varepsilon_{xy}(\tau)$ [138, 137, 38, 93]. This noise has zero mean $\langle\mathrm{d}\varepsilon_{xy}(\tau)\rangle=0$ and the “white noise” property $\langle\mathrm{d}\varepsilon_{xy}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle_{x_{\tau}=k}=\delta_{ix}\delta_{jy}\delta_{ik}\delta(\tau-\tau^{\prime})r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$, where the expectation is conditioned on $x_{\tau}=k$. The Kronecker $\delta$ show that the noise increments on different edges are independent and only diagonal elements survive, e.g., $\langle\mathrm{d}\varepsilon_{ij}(\tau)\mathrm{d}\varepsilon_{ij}(\tau^{\prime})\rangle_{x_{\tau}=k}=\delta_{ik}\delta(\tau-\tau^{\prime})r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}$. Moreover, since $|\mathrm{d}\varepsilon_{ij}|\sim\sqrt{\mathrm{d}\tau}$, higher moments scale as $|\mathrm{d}\varepsilon_{ij}|^{n}\sim(\mathrm{d}\tau)^{n/2}$, so that all higher moments and cumulants with $n\geq 3$ vanish for small $\mathrm{d}\tau$. Note that the Langevin Eq. (2) has already been introduced in [93].
We prove the central time-noise correlation lemma with $\mathrm{d}\tau_{i}(\tau)={\color[rgb]{0,0,0}\delta_{x_{\tau}i}}\mathrm{d}\tau$,
@@
$=\delta_{xk}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau),$
@@
where $P(x,\tau^{\prime}|i,\tau)$ is the transition probability (i.e., the “propagator”) from state $i$ at $\tau$ to state $x$ at $\tau^{\prime}\geq\tau$ and $p_{i}(\tau)$ is the probability to find the system in state $i$ at time $\tau$. Note that the averages are “pinned”, e.g., $\langle\cdot\rangle^{x_{\tau^{\prime}}=x}_{x_{\tau}=i}$ is the average conditioned on $x_{\tau}=i$} and $x_{\tau^{\prime}}=x$} at times $\tau\leq\tau^{\prime}$. The above correlations were shown in [93] using a different approach.
@@
$J_{t}^{\mathrm{c}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\circ\mathrm{d}\bm{x}_{\tau},$ (Stratonovich)
@@
for continuous-space and MJP, respectively, where the former is a Stratonovich integral over some vector-valued $\bm{U}(\bm{x}_{\tau},\tau)$ and the $T$ denotes the transpose. Recall that the Stratonovich integral is a Riemann–Stieltjes-type integral defined as the limit of of functions evaluated at midpoints, e.g., $\bm{f}(\bm{x}_{\tau})^{T}\circ\mathrm{d}\bm{x}_{\tau}=\bm{f}(\frac{\bm{x}_{\tau+\mathrm{d}\tau}+\bm{x}_{\tau}}{2})^{T}(\bm{x}_{\tau+\mathrm{d}\tau}-\bm{x}_{\tau})$. To give an intuition for $\bm{U}$, consider the example $\bm{U}(\bm{x}_{\tau},\tau)]_{i}={\color[rgb]{0,0,0}\mathbb{1}}_{\bm{x}_{\tau}\in\{-h/2,h/2\}^{d}}$ the indicator valued vector being non-zero if $\bm{x}_{\tau}$ is in the hypercube with side length $h$, then the current “counts up” the displacement of the trajectory $\{\bm{x}_{\tau}\}_{0\leq\tau\leq t}$ in that hypercube. Note the superscript c throughout this manuscript refers to continuous space quantities and we write superscript (c) as a placeholder when the corresponding relation is valid in both continuous and discrete space systems. For MJP the transitions are weighted by a time-dependent antisymmetric transition weight $\bm{\kappa}(\tau)=-\bm{\kappa}(\tau)$ which ensures a sign change in the current when reversing time [144]. Hence, we identify Stratonovich integrals for MJP as integrals over traces of a two-point function and the jump increments. The above currents naturally decompose into a stochastic and “usual” integral [38], $J^{(\mathrm{c})}_{t}=J_{t}^{\mathrm{(c),I}}+J_{t}^{\mathrm{(c),II}}$, where the stochastic part is
$J_{t}^{\mathrm{c,I}}$ $=\int_{\tau=0}^{\tau=t}\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$ (Itô)
@@
respectively, while the usual” part reads $J_{t}^{\mathrm{(c),II}}=\int_{0}^{t}\mathcal{U}_{\tau}^{(c)}\mathrm{d}\tau$ where $\mathcal{U}_{\tau}^{c}\equiv\bm{U}(\bm{x}_{\tau},\tau)^{T}\bm{F}(\bm{x}_{\tau})+\nabla\cdot\left[\bm{D}\bm{U}(\bm{x}_{\tau},\tau)\right]$, $\bm{D}=\bm{\sigma}\bm{\sigma}^{T}/2$ is the diffusion matrix, and $\mathcal{U}_{\tau}\equiv\mathrm{Tr}[\bm{\kappa}(\tau)^{T}\bm{R}(x_{\tau})]$. Thus, we identify in the second line of Eq. (6) the Ito-type integral for observables of MJP by analogy to diffusions.
@@
$\rho_{t}$ $=\int_{\tau=0}^{\tau=t}{\color[rgb]{0,0,0}V_{\tau}\mathrm{d}\tau},$ (7)
for continuous and discrete dynamics, respectively, where $V(\bm{x}_{\tau},\tau)$ and $V_{\tau}\equiv V(x_{\tau},\tau)=\sum_{k}\delta_{x_{\tau}k}V_{k}(\tau)$ are scalar-valued functions.
@@
$\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{\alpha},{\color[rgb]{0,0,0}G}^{\beta}}[\cdot]$ $=\int_{0}^{t}\mathrm{d}\tau\int_{\tau}^{t}\mathrm{d}\tau^{\prime}\sum_{x,y}{\color[rgb]{0,0,0}G}_{xy}^{\alpha}\sum_{i,j}{\color[rgb]{0,0,0}G}^{\beta}_{ij}[\cdot],$ (8)
for continuous and discrete dynamics respectively. The operator in the first line of Eq (8) is already known [36, 35]. For $X_{t}^{\mathrm{(c)}}\in\{J_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}\}$ and $Y_{t}^{\mathrm{(c)}}\in\{J_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}},\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}\}$, $k,l\in\{1,2\}$, we find
@@
$\mathrm{cov}(X_{t},Y_{t})$ $=\hat{{I}}^{t}_{{\color[rgb]{0,0,0}G}^{X},{\color[rgb]{0,0,0}G}^{Y}}[{\Xi}^{m}_{ijxy}]-\langle X_{t}\rangle\langle Y_{t}\rangle,$ (9)
where $X_{t}^{\mathrm{(c)}},Y^{\mathrm{(c)}}_{t}$ are the appropriate pathwise observables and ${\color[rgb]{0,0,0}G}^{X},{\color[rgb]{0,0,0}G}^{Y}$ functions that define them ( e.g., $\bm{\kappa}$ in the case of currents for MJP). Moreover, $m\in\{1,2,3\}$ indicates density-density ($m=1$), current-density ($m=2$), and current-current ($m=3$) covariances. For the density covariance ($m=1$), $X_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}k}}^{\mathrm{(c)}}$ and $Y_{t}^{\mathrm{(c)}}=\rho_{t,{\color[rgb]{0,0,0}l}}^{\mathrm{(c)}}$, we get
${\Xi}_{1}^{\bm{z}\bm{z}^{\prime}}$ $=\langle{\color[rgb]{0,0,0}\mathbb{1}}\rangle_{\bm{x}_{\tau}=\bm{z}}^{\bm{x}_{\tau^{\prime}}=\bm{z}^{\prime}}+\langle{\color[rgb]{0,0,0}\mathbb{1}}\rangle_{\bm{x}_{\tau}=\bm{z}^{\prime}}^{\bm{x}_{\tau^{\prime}}=\bm{z}},$ (10)
@@
$\mathrm{cov}({\rho}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{V^{k}\delta,V^{l}\delta}\left[P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})+P_{\mathrm{s}}(x,\tau^{\prime};i,\tau)-p_{x}^{\mathrm{s}}p_{i}^{\mathrm{s}}\right]$
$\mathrm{cov}({J}_{t,{\color[rgb]{0,0,0}k}},{\rho}_{t,{\color[rgb]{0,0,0}l}})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\hat{j}_{xy}P(x,t^{\prime}|i)p_{i}^{\mathrm{s}}+\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}p_{i}^{\mathrm{s}}\right],$
$\mathrm{cov}(J_{t,{\color[rgb]{0,0,0}k}},{J}_{t,{\color[rgb]{0,0,0}l}})=$ $t\sum_{x,y}\kappa_{xy}^{k}\kappa_{xy}^{l}r_{xy}p_{x}^{\mathrm{s}}+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\hat{j}_{xy}\hat{j}_{ji}^{\ddagger}P(x,t^{\prime}|j)p_{j}^{\mathrm{s}}+\hat{j}_{ij}\hat{j}_{yx}^{\ddagger}P(i,t^{\prime}|y)p_{y}^{\mathrm{s}}-J^{\mathrm{s}}_{xy}J^{\mathrm{s}}_{ij}\right],$ (13)
where $V^{\alpha/\beta}\delta$ is a shorthand notation for $V^{\alpha/\beta}_{i}\delta_{ij}$, $J^{\mathrm{s}}_{xy}$ is the local steady-state current between $x$ and $y$, and $\hat{j}_{xy}^{\ddagger}$ is the so-called dual-reverse current operator (reversed steady-state current) [37, 36, 35] which for MJP reads $\hat{j}_{xy}=r_{xy}\to\hat{j}_{xy}^{\ddagger}=r_{yx}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$ [144]. In Eq. (13), the joint stationary probabilities for state $x$ at $\tau$ and $i$ at $\tau^{\prime}$ are denoted $P_{\mathrm{s}}(x,\tau;i,\tau^{\prime})$, while the marginal stationary distributions to be in a state $x$ is $p^{\mathrm{s}}_{x}$. Additionally, we use the fact that the propagator for time-homogeneous systems is time-translation invariant $P(i,\tau^{\prime}|x,\tau)=P(i,\tau^{\prime}-\tau|x)$. The results for continuous-space diffusion are given in Refs. [36, 35]. For stationary initial conditions and with time-independent state function and transition weights, Eqs. (13) have the form of generalized Green-Kubo-like relations, in the sense that they relate (co)variances of observables to time integrals of (generalized) correlation functions of density- and current-observables (see Refs. [37, 36] for a discussion of diffusion processes). Notably, Eqs. (13) have a striking similarity to the continuous space expressions [36, 37]. Later we show how to to generalize Eqs. (13) to include time-dependent transition weights and state functions, as well as transient dynamics.
@@
where $c(t):{\color[rgb]{0,0,0}\mathbb{R}}\to{\color[rgb]{0,0,0}\mathbb{R}}$ is a deterministic weight function for the density and $\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}$ is the discrete (continuous) space entropy production [143]
$\Delta S_{\mathrm{tot}}^{\mathrm{c}}(t)$ $=\int_{0}^{t}\mathrm{d}\tau\int\mathrm{d}\bm{x}\frac{\bm{j}(\bm{x},\tau)^{T}\bm{D}^{-1}\bm{j}(\bm{x},\tau)}{P(\bm{x},\tau)},$ (15)
$\Delta S_{\mathrm{tot}}(t)$ $=\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}r_{ij}p_{i}(\tau)\log\frac{r_{ij}p_{i}(\tau)}{r_{ji}p_{j}(\tau)}.$ (16)
Here, $\bm{j}(\bm{x},\tau)=(\bm{F}(\bm{x})-\bm{D}\nabla)P(\bm{x},\tau)$ is the probability current entering the Fokker Planck equation. In analogy with [38] we introduced the discrete space modified time-integrated density and current $\tilde{\rho}_{t}=\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}{\color[rgb]{0,0,0}V_{\tau}\mathrm{d}{\tau}}$ and $\tilde{J}_{t}=\int_{\tau=0}^{\tau=t}\tau\mathrm{Tr}\left[(\partial_{\tau}\bm{\kappa}(\tau))^{T}\mathrm{d}\bm{n}(\tau)\right]$ that account for the dependence of $V_{i}(\tau)$ and $\bm{\kappa}(\tau)$ on $\tau$. To compare, the continuous space modified current and density read $\tilde{J}_{t}^{\mathrm{c}}=\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}\bm{U}(\bm{x}_{\tau},\tau)^{T}\circ\mathrm{d}\bm{x}_{\tau}$ and $\tilde{\rho}_{t}^{\mathrm{c}}=\int_{0}^{t}\tau\partial_{\tau}V(\bm{x}_{\tau},\tau)\mathrm{d}\tau$, respectively. The transient correlation TUR for MJP in Eq. (14) has not yet been proven in this generality since previous results were restricted to steady states [32] or did not consider correlations [89].
Setting $c(t)=0$ we recover the transient TUR. The CTUR (in contrast to the TUR) allows to turn the Cauchy-Schwarz inequality into an equality for all times given a suitable choice of $V_{\tau}$ and $\bm{\kappa}(\tau)$. This choice, however, requires explicit and detailed information about the system and is therefore not applicable in most realistic applications. We can nevertheless optimize Eq. (14) with respect to $c(t)$ for given $V_{\tau}$ and $\bm{\kappa}(\tau)$ to find